OpenStudy (anonymous):

Solve the following equation ln(2x+3)+ln(x-6)=2lnx

OpenStudy (anonymous):

simplify using rules of logs then take e^ both sides to get rid of the lns

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

What did you get as your result so far after doing what Peter told you?

OpenStudy (anonymous):

@WilliamF You there?

OpenStudy (anonymous):

@Azteck Yes, sorry just been a while since I've used log, I'll let you know when I get an answer

OpenStudy (anonymous):

Okay, no worries.

OpenStudy (anonymous):

Just for your benefit, you need to use the quadratic formula afterwards to find x, since you won't be able to factorise it.

OpenStudy (anonymous):

@kyusakazaki Here's some tips for you when helping people on OpenStudy: 1. Please don't give answers away without letting the person know what you're trying to do. 2. Actually guide the person step by step. 3. This is not really a tip, but "ln" is not log to the base 10. It seems you haven't gotten that far ahead in maths to know that "ln" is log to the base "e".

OpenStudy (anonymous):

thanks aztek i made a bad mistake

OpenStudy (anonymous):

No worries man.

OpenStudy (anonymous):

so using rules of logs I get \[\ln \frac{ 2x^{2}-9x+12 }{ x ^{2}}\] Can't you just cancel out the x^2 then find an answer of x? Although I have a feeling you can't do this

OpenStudy (hexagon001):

where did you get the denominator from?

OpenStudy (anonymous):

Only simplify the LHS. Don't touch anything on the RHS first.

OpenStudy (anonymous):

Using log laws: \[ln(a)+ln(b)=ln(ab)\] \[ln(2x+3)+ln(x-6)=ln(2x+3)(x-6)\] Do you follow so far?

OpenStudy (anonymous):

@WilliamF You there?

OpenStudy (anonymous):

Yes I'm here and follow that part

OpenStudy (anonymous):

So that's the LHS done.

OpenStudy (anonymous):

Now the RHS. Using the log law: \[xlna=lna^{x}\] The RHS will become: \[2lnx=lnx^2\]

OpenStudy (anonymous):

Do you follow this part as well?

OpenStudy (anonymous):

yep, that is where I got the denominator btw

OpenStudy (anonymous):

Now we have this: \[\large ln(2x+3)(x-6)=lnx^2\] WE can just get rid of the ln. and we will then have: \[\large (2x+3)(x-6)=x^2\]

OpenStudy (anonymous):

Do you get that? We just e^ both sides.

OpenStudy (anonymous):

We don't need to bring the lnx^2 to the LHS.

OpenStudy (anonymous):

It's much easier to do that.

OpenStudy (anonymous):

If you get that part, then you can continue on expanding the LHS and moving everything to one side. THen you can use the quadratic formula to solve for x. Could you please give me your answer after you've done those steps I've listed above. Cheers.

OpenStudy (anonymous):

I got \[x ^{2}-9x+72\] for my final answer and for the values of x I got x=10.68 and x=-1.68

OpenStudy (anonymous):

where did you get 72 from?

OpenStudy (anonymous):

\[3\times -8\neq72\]

OpenStudy (anonymous):

Sorry I meant 18, not sure how that got there

OpenStudy (anonymous):

Good. Excellent.

OpenStudy (anonymous):

Now, you have to check each answer to see if it works with the original equation.

OpenStudy (anonymous):

Could you do that please and see if the negative answer works in the equation.

OpenStudy (anonymous):

Use a calculator.

OpenStudy (anonymous):

If it doesn't then, the negative answer is not a solution. Therefore you would only have one, since I'm sure the positive answer would work.

OpenStudy (anonymous):

Yeah I get a math error for the negative answer

OpenStudy (anonymous):

So therefore, it doesn't work. You will thus have one solution.

OpenStudy (anonymous):

So you would put an arrow to the negative answer and write no solution. Then on the next line you write, therefore x=10.68

OpenStudy (anonymous):

Or something similar to that.

OpenStudy (anonymous):

Ok that makes sense, thanks for the help it is much appreciated

OpenStudy (anonymous):

Is there a way to give a medal or something, I'm new here.

OpenStudy (anonymous):

No worries. IF you need any clarification, you should ask your teacher for further information. Yes, there is, on he right hand side, it says best response. You can choose the person that helped you the most or gave you a good answer.

OpenStudy (anonymous):

And I almost forgot that you were new here: Welcome to OpenStudy, feel free to ask anymore questions on here, if you feel you're stuck on something. Have fun and enjoy using this website properly and efficiently.

OpenStudy (anonymous):

on the*